OpenStudy (anonymous):

What are the root(s) of the function f(x) = 2x 2 - 7x - 15? x= 5 x= -, 5 x= -2, 15 x= 1,

OpenStudy (anonymous):

put f(x)=0 2x^2-10x+3x-15=0 2x(x-5)+3(x-5)=0 x=-3/2 and x=5